What is the common difference and generale rule for generating the terms of the sequence for A1=8,A17=88 and how to you resolve this problem?
Answer by prasadmlm(2)

here we have to calculate common difference which means given progression is arthemetic progression.
given first term a1=8
seventeen th term a7=88
but a17=a1+16d where a1 means first term and d means common difference
therfore a1+16d=88
=>8+16d=88
=>16d=88-8
=>16d=80
=>d=80/16
=>d=5
therfore common difference d is 5.
general term tn=a1+(n-1)d
=8+(n-1)5
=8+5n-5
=5n+3.
and progression is 8,13,18,23,28,------------------